What is the surface area of a shoe box that is 12 inches in length 5 inches in width and 5 inches in height
Answer by richwmiller(17219)

Ignoring the fact the the shoe box lid is slightly larger than the bottom of the Box.
The box has 4 sides of the same size 4*(12*5)=240 and the two ends of the same size 2*(5*5)=50
240+50=290